OpenStudy (anonymous):

Translate the sentence into an equation. Four times a number is 24. A. 4n = 24 B. 4 + n = 24 C. 4/n=24 D. n = 4 ∙ 24

OpenStudy (cj49):

let the number be n. four times a number means 4 multiplied by n is 24 so now wat wud be the equation?

OpenStudy (anonymous):

A

OpenStudy (cj49):

its correct..

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Which equation matches the sentence? Column A Column B 1. Seven more than 2 times a number is 23. 2. The difference of 2 times a number and 7 is 23. 3. The sum of 7 times a number and 2 is 23. 4. Seven less than 2 times a number is 23. A. 7n + 2 = 23 B. 2n – 7 = 23 C. 2n + 7 = 23

OpenStudy (anonymous):

1 2 3 4 = column A A B C = column B

OpenStudy (cj49):

more means addition and y times a number means y(n) difference means minus

OpenStudy (cj49):

now wud u be able to do it from here...

OpenStudy (anonymous):

probably not but i can try

OpenStudy (cj49):

yes pls....ill help u out if u wnt understand any equation..

OpenStudy (anonymous):

i figured it out

OpenStudy (anonymous):

1=c 2=b 3=a 4=b

OpenStudy (cj49):

correct...

OpenStudy (anonymous):

5. Charlie has 8 bills in his wallet. Five of the bills are $10 bills, and the other bills are $5 bills. What is the total value of the 8 bills in Charlie's wallet? The total value of the 8 bills is $ .